# NaviAgent: Bilevel Planning on Tool Dependency Graphs for Function Calling

### v1

@@ B.3 Graph Search Algorithm
We adapt the Alpha-Beta pruning algorithm for backward search over the tool dependency graph $\mathcal{G}=(V,E,W)$, parameterized by a quintuple $(\alpha,\beta,\mathcal{H},\mathcal{D},\mathcal{C})$, where $\alpha\in\mathbb{R}^{+}$ (initialized as $\alpha_{0}=0.4$) is the lower-bound threshold for acceptable path scores, and $\beta\in\mathbb{R}^{+}$ (with $\beta_{0}=0.9$) is the upper-bound for candidate evaluation. The dynamic threshold function $\mathcal{H}(d)=\max(0.3,0.5\times 0.9^{d})$ applies exponential decay to balance search depth $d$ and semantic relevance. The depth attenuation factor $\mathcal{D}(d)=1/(1+\sqrt{d})$ penalizes longer paths. The connectivity constraint $\mathcal{C}(u,v_{t})=\mathrm{PathLength}(u,v_{t})\leq 5$ ensures that generated subgraphs remain compact, where $v_{t}$ denotes the target node (either an API node or a parameter node). The parametric scoring function is defined as:
@@
The algorithm is parameterized by a sextuple $(\mathcal{T}_{0},\eta,\mathcal{P},d_{\max},\mathcal{M}_{\theta},\mathcal{F}_{\omega})$ (see Algorithm 2), where $\mathcal{T}_{0}=300$ is the initial temperature that determines the probability of accepting suboptimal solutions and balances exploration and exploitation, $\eta=0.7$ is the cooling rate that controls the annealing schedule $\mathcal{T}_{k+1}=\eta^{1+k/5}\mathcal{T}_{k}$, $\mathcal{P}=40$ is the population size, $d_{\max}=4$ is the maximum search depth, and $\mathcal{M}_{\theta}$ is a temperature-sensitive mutation operator with adaptive intensity $\theta=\lfloor\mathcal{T}/100\rfloor$. Candidate solutions are evaluated using a composite fitness function:
@@
Dependency graph $G$, target API set $A$, initial temperature $\mathcal{T}_{0}=300$, cooling rate $\eta=0.7$, population size $\mathcal{P}=40$, maximum search depth $d_{max}=4$
@@
6:     Generate initial population $Pop$ (size $\mathcal{P}$) for API $a$
@@
         Evaluate fitness $\mathcal{F}_{\omega}$ for each chromosome in $Pop$
@@
         Update population $Pop$ with offspring
@@
     Select best chromosome from $Pop$ based on $\mathcal{F}_{\omega}$
@@ Appendix C Cases
The following three cases exemplify the Decider-Navigator collaborative mechanism through four core actions executed by the Decider: 1) Direct Response: resolves user queries using pre-trained knowledge. 2) Intent Clarification: initiates interactive dialogue to disambiguate vague requests. 3) Tool Retrieval: collaborates with the Navigator module to generate a pruned tool dependency subgraph, which is returned as executable toolchains. 4) Tool Call: using above dependency subgraph, Executes API with parameter validation and state monitoring. This design achieves centralized decision control through the Decider’s orchestration authority while enabling dynamic resource optimization via the Navigator’s graph-based toolchain generation, ensuring both operational efficiency and system integrity in complex task environments.
@@ C.1 Case 1
    "analysis": "1. Common sense judgment: Gastroenteritis is a medical
    condition within general knowledge. 2. No API required for definition
    explanation. 3. Parameter verification not applicable",
@@
    "intent": "1. Provide medical definition of gastroenteritis 2. User’s
    explicit request for condition explanation",
@@
    "answer": "Gastroenteritis is inflammation of the stomach and intestines,
    typically resulting from bacterial toxins or viral infection. Common
@@
    ACTION: Tool Retrieval
@@
    "analysis": "1. Health data recording requires specialized API 2. No
    matching API in candidate list 3. Critical parameters provided:
    user_id=12345, blood_pressure=130/80, heart_rate=90, timestamp=2025-03-05",
@@
    "intent": "1. Store user health metrics 2. Provided parameters: blood
    pressure, heart rate, user ID, timestamp",
@@
    "recall_description": "HealthDataAPI(description: Records user health
@@
    blood_pressure:string/systolic & diastolic, heart_rate:integer/bpm,
    timestamp:datetime/measurement time; output: record_id:string/new entry
    identifier)",
    "answer": "I need to retrieve a health data recording API to complete this
    action. Should I proceed with API retrieval?"
    Then, the BGE model retrieves three candidate APIs (api-RecordHealthData, api-QueryHealthData, api-QueryRegistration) based on the recall description above, after which the Navigator module is invoked to acquire the corresponding call chains for these APIs (see Figure 8).
@@
In this round, the Decider module selects and executes the optimal API chain, with api-RecordHealthData designated as the target API.
ACTION: Tool Call
@@
ACTION: Tool Call
@@ 2. Health Data Record Confirmation:
    All parameters validated successfully through api-RecordHealthData. No data
    integrity issues detected.
@@ C.2 Case 2
      ’analysis’: ’1. The specific problem addressed by this action is to determine
      if Beijing is located within China. 2. The user input directly asks about the
      geographical location of Beijing. 3. This is a common sense question and does
      not require an API call. 4. No parameters are needed for this decision.’,
@@
    ACTION: Tool Retrieval
@@
    ’intent’: ’What is the current weather forecast for the location at latitude
    40.7128 and longitude -74.0060 in metric units?’,
    ’analysis’: ’1. The specific problem addressed by this action is to retrieve
    the current weather forecast for a specific location using latitude and
    longitude coordinates. 2. This is not a common sense question and requires an
    external API call. 3. The API matching phase will prioritize APIs that can
    provide weather forecasts based on geographic coordinates. 4. Required
    parameters are latitude and longitude, which are provided in the user input.’,
@@
    ’recall_description’: ’WeatherAPI(description: Retrieve current weather
    conditions by geographic coordinates, input: latitude:float/latitude
    coordinate; longitude:float/longitude coordinate; output:
    temperature:float/current temperature; humidity:float/current humidity;
@@
    _weather) based on the recall description above, after which the Navigator module is invoked to acquire the corresponding call chains for these APIs (see Figure 9).
@@
In this round, the Decider module selects and executes the optimal API chain, with api-by_geographic
@@
ACTION: Tool Call
@@
  ’analysis’: "1. The specific problem addressed by this action is to retrieve
  the current weather forecast for a specific location using the provided
  latitude and longitude. 2. The user input directly provides the necessary
  parameters (latitude and longitude) and specifies the need for metric units.
  3. The target API ’api-current_weather_data_of_a_location_for_weather’ has
  already been successfully called and its result is available. 4. The required
  parameters (lat and lon) are complete and the API call has been confirmed
@@ 3. The target API ’api-current_weather_data_of_a_location_for_weather’ has
  ’answer’: ’The current weather in New York is clear sky with a temperature of
  11.0C, feels like 8.0C, and the wind speed is 5.1 m/s from the west-
  southwest.’}
@@
Beijing is in China. The current weather in New York is clear sky with a
temperature of 11.0C, feels ike 8.0C, and the wind speed is 5.1 m/s from the west-
southwest.
@@ C.3 Case 3
    "analysis": "1. Weather data needs to be queried in real time -> not common
@@
    "answer": "Which city do you want to query tomorrow’s weather?"
@@
    ACTION: Tool Retrieval
@@
      ’intent’: ’1. Query the weather forecast for tomorrow in Beijing 2. Extract
      location: Beijing and time: tomorrow from user input’,
      ’analysis’: ’1. The specific problem addressed by this action is to retrieve
      the weather forecast for tomorrow in Beijing. 2. The user input directly
      provides the location (Beijing) and the time frame (tomorrow). 3. Since no
      candidate APIs are available, the system needs to retrieve an appropriate API
      for weather forecasting. 4. There are no parameters provided by the user that
      can be used directly with an API, so the system must retrieve an API that can
@@
      ’recall_description’: ’WeatherForecastAPI(description: Retrieve weather
      forecast for a given location and date, input: location:string/city name;
      date:date/forecast date; output: weather:string/weather condition,
      temperature:float/forecast temperature)’}
@@
    _weather) based on the recall description above, after which the Navigator module is invoked to acquire the corresponding call chains for these APIs (see Figure 10).
@@
In this round, the Decider module selects and executes the optimal API chain, with api-getweatherforecast_for_apjoy_weather_forecast designated as the target API.
ACTION: Tool Call
@@
  The weather forecast for tomorrow in Beijing is partly cloudy with a
  temperature of 25C and humidity of 60%.
@@ 1. Core Requirements:
      Must explicitly contain initial parameters for leaf-node APIs
      Implicitly requires chained API calls from leaf to root node
      Root node API’s output directly resolves the user’s problem
@@ 2. Dependency Chain Rules:
      The output of every leaf-node API must be utilized in downstream
@@ 3. Parameter Constraints:
   - Root API output must contain realistic values matching its schema
@@ 5. Response Structure:
  "query": "<Real-world scenario requiring sequential API calls>",
@@
        "<param>": "<value explicitly stated in user query or previous
                   API output>"
@@
        "data": {"<field>": "<realistic resolution to query>"}
@@ E.3.1 NaviAgent Inference Prompts
You are an intelligent API coordination system. Respond strictly according
to the following rules:
@@
   - Decompose compound requests into independent ordered sub-intents
      Sequential dependencies first, Must execute in declared order
@@
       - Create self-contained decision unit, action must implement full
@@
       - Each action’s analysis focuses only on its own intent
@@ 2. Match intent with API list:
        - Validate Observation in user input to confirm target API success:
@@
            a) Complete parameters -> action=call_api (execute based on 3.1
            dependency resolution)
@@
       a) Can be completed via dependent APIs -> Execute Rule 3.1
       b) Use Retrieval APIs resolve parameter deficiencies in API
@@
    a) Check required parameters of target API, first call dependent APIs.
    b) For each missing parameter, select APIs from dependencies not marked
@@
    c) If an input parameter of an API is unavailable, use retrieve_api to
@@
  "intent": "1. <precisely describe the specific problem addressed by the
        current action>
             2. <extract data segments directly related to the subtask from
        user input>",
@@
        1.Explicitly state the specific decision-making sub-intent
@@
        "When action=retrieve_api: Use ’APIName(description: API
        functionality, input: param:type/description; output:
        param:type/description)’ format with only core parameters (e.g.,
        ’StockAPI(description: Query stock price by symbol, input:
        symbol:string/stock symbol; output: price:float/current price)’)",
  "answer": "When actionin[direct_answer,clarify_intent]: Natural language
        response (interrogative sentences required)"
@@
- JSON array items MUST be sorted by dependency_rank in ascending order
@@
[OR]Target API must be executed successfully, as shown in the status
@@ 1. Parameter names must strictly match API documentation
2. The ’answer’ field for clarify_intent must contain question words
@@ 4. When action in [retrieve_api]:
    - The recall_description field serves exclusively as an API retrieval
      identifier from predefined repositories.
    - parameter descriptions must distinguish between input and output
      parameters, retaining only essential parameters
    - Each recall_description can only recall one api,multiple APIs require
@@ 5. APIs absent from Candidate APIs MUST NOT be invented
6. When action=call_api is permitted only when candidate APIs exist and the
   target_api is present in the candidate APIs.
7. The "action" field must be strictly limited to one of the following four
   predefined operation types: call_api, direct_answer,retrieve_api or
   clarify_intent.

### v2

@@ B.2 Graph Search Algorithm
This algorithm knuth1975analysis is adapted for backward search over the tool dependency graph $\mathcal{G}=(V,E,W)$, parameterized by a quintuple $(\alpha,\beta,\mathcal{H},\mathcal{D},\mathcal{C})$, where $\alpha\in\mathbb{R}^{+}$ (initialized as $\alpha_{0}=0.4$) is the lower-bound threshold for acceptable path scores, and $\beta\in\mathbb{R}^{+}$ (with $\beta_{0}=0.9$) is the upper-bound for candidate evaluation. The dynamic threshold function $\mathcal{H}(d)=\max(0.3,0.5\times 0.9^{d})$ applies exponential decay to balance search depth $d$ and semantic relevance. The depth attenuation factor $\mathcal{D}(d)=1/(1+\sqrt{d})$ penalizes longer paths. The connectivity constraint $\mathcal{C}(u,v_{t})=\mathrm{PathLength}(u,v_{t})\leq 5$ ensures that generated subgraphs remain compact, where $v_{t}$ denotes the target node (either an API node or a parameter node). The parametric scoring function is defined as:
@@
Our hybrid heuristic search algorithm combines simulated annealing kirkpatrick1983optimization and genetic algorithm strategies shapiro1999genetic . It is parameterized by a sextuple $(\mathcal{T}_{0},\eta,\mathcal{P},d_{\max},\mathcal{M}_{\theta},\mathcal{F}_{\omega})$ (see Algorithm 2), where $\mathcal{T}_{0}=200$ is the initial temperature that determines the probability of accepting suboptimal solutions and balances exploration and exploitation, $\eta=0.7$ is the cooling rate that controls the annealing schedule $\mathcal{T}_{k+1}=\eta^{1+k/5}\mathcal{T}_{k}$, $\mathcal{P}=20$ is the population size, $d_{\max}=4$ is the maximum search depth, and $\mathcal{M}_{\theta}$ is a temperature-sensitive mutation operator with adaptive intensity $\theta=\lfloor\mathcal{T}/100\rfloor$. Candidate solutions are evaluated using a composite fitness function:
@@
We parallelize the subgraph search for different target APIs in Algorithm 2. This approach processes the population evolution tasks independently and concurrently, thereby eliminating the computational bottleneck of the original algorithm’s serial loops.
@@
Dependency graph $G$, target API set AA, initial temperature $\mathcal{T}_{0}=200$, cooling rate $\eta=0.7$, population size $\mathcal{P}=20$, maximum search depth $d_{max}=4$
@@
6:  Generate initial population PopPop (size $\mathcal{P}$) for API $a$
@@
   Evaluate fitness $\mathcal{F}_{\omega}$ for each chromosome in PopPop
@@
   Update population PopPop with offspring
@@
  Select best chromosome from PopPop based on $\mathcal{F}_{\omega}$
@@ Appendix C Cases
The following three cases exemplify the bilevel planning mechanism through four core actions: 1) Direct Response: resolves user queries using pre-trained knowledge. 2) Intent Clarification: initiates interactive dialogue to disambiguate vague requests. 3) ToolChain Retrieval: works with the TWNM to construct a pruned tool dependency subgraph, which is then returned as an executable toolchain. 4) Tool Execution: executes the required APIs based on the dependency subgraph, with parameter validation and state monitoring. This design achieves centralized decision control through the agent’s orchestration authority while enabling dynamic resource optimization via the TWNM’s graph-based toolchain generation, ensuring both efficiency and robustness of the our framework in complex task environments.
@@ C.1 Case 1
    "analysis": "1. Common sense judgment: Gastroenteritis is a
    medical condition within general knowledge. 2. No API
    required for definition explanation. 3. Parameter
    verification not applicable",
@@
    "intent": "1. Provide medical definition of gastroenteritis
    2. User’s explicit request for condition explanation",
@@
    "answer": "Gastroenteritis is inflammation of the stomach
    and intestines,
    typically resulting from bacterial toxins or viral
    infection. Common
@@
    ACTION: TooChain Retrieval
@@
    "analysis": "1. Health data recording requires specialized
    API 2. No matching API in candidate list 3. Critical
    parameters provided:
    user_id=12345, blood_pressure=130/80, heart_rate=90,
    timestamp=2025-03-05",
@@
    "intent": "1. Store user health metrics 2. Provided
    parameters: blood pressure, heart rate, user ID, timestamp",
@@
    "recall_description": "HealthDataAPI(description: Records
    user health
@@
    blood_pressure:string/systolic & diastolic,
    heart_rate:integer/bpm,
    timestamp:datetime/measurement time; output:
    record_id:string/new entry identifier)",
    "answer": "I need to retrieve a health data recording API to
    complete this action. Should I proceed with API retrieval?"
    Then, the BGE model retrieves two candidate APIs (api-RecordHealthData, api-QueryRegistration) based on the recall description above, after which the TWNM is invoked to acquire the corresponding call chains for these APIs (see Figure 8).
@@
In this round, the agent selects and executes the optimal API chain, with api-RecordHealthData designated as the target API.
ACTION: Tool Execution
@@
ACTION: Tool Execution
@@ 2. Health Data Record Confirmation:
    All parameters validated successfully through api-
    RecordHealthData. No data integrity issues detected.
@@ C.2 Case 2
      ’analysis’: ’1. The specific problem addressed by this
      action is to determine if Beijing is located within China.
      2. The user input directly asks about the
      geographical location of Beijing. 3. This is a common
      sense question and does not require an API call. 4. No
      parameters are needed for this decision.’,
@@ 2. The user input directly asks about the
    ACTION: TooChain Retrieval
@@
    ’intent’: ’What is the current weather forecast for the
    location at latitude 40.7128 and longitude -74.0060 in
    metric units?’,
    ’analysis’: ’1. The specific problem addressed by this
    action is to retrieve the current weather forecast for a
    specific location using latitude and longitude coordinates.
    2. This is not a common sense question and requires an
    external API call. 3. The API matching phase will prioritize
    APIs that can provide weather forecasts based on geographic
    coordinates. 4. Required parameters are latitude and
    longitude, which are provided in the user input.’,
@@ 2. This is not a common sense question and requires an
    ’recall_description’: ’WeatherAPI(description: Retrieve
    current weather conditions by geographic coordinates, input:
    latitude:float/latitude coordinate;
    longitude:float/longitude coordinate; output:
    temperature:float/current temperature;
    humidity:float/current humidity;
@@
    _weather) based on the recall description above, after which the TWNM is invoked to acquire the corresponding call chains for these APIs (see Figure 9).
@@
In this round, the agent selects and executes the optimal API chain, with api-by_geographic
@@
ACTION: Tool Execution
@@
  ’analysis’: "1. The specific problem addressed by this
  action is to retrieve the current weather forecast for a
  specific location using the provided latitude and
  longitude. 2. The user input directly provides the
  necessary parameters (latitude and longitude) and
  specifies the need for metric units.
  3. The target API ’api-
  current_weather_data_of_a_location_for_weather’ has
  already been successfully called and its result is
  available. 4. The required parameters (lat and lon) are
  complete and the API call has been confirmed
@@ 3. The target API ’api-
  ’answer’: ’The current weather in New York is clear sky
  with a temperature of 11.0C, feels like 8.0C, and the
  wind speed is 5.1 m/s from the west-southwest.’}
@@
Beijing is in China. The current weather in New York is
clear sky with a temperature of 11.0C, feels ike 8.0C, and
the wind speed is 5.1 m/s from the west-southwest.
@@ C.3 Case 3
    "analysis": "1. Weather data needs to be queried in real
@@
    "answer": "Which city do you want to query tomorrow’s
    weather?"
@@
    ACTION: TooChain Retrieval
@@
      ’intent’: ’1. Query the weather forecast for tomorrow in
      Beijing 2. Extract location: Beijing and time: tomorrow
      from user input’,
      ’analysis’: ’1. The specific problem addressed by this
      action is to retrieve
      the weather forecast for tomorrow in Beijing. 2. The user
      input directly provides the location (Beijing) and the
      time frame (tomorrow). 3. Since no candidate APIs are
      available, the system needs to retrieve an appropriate API
      for weather forecasting. 4. There are no parameters
      provided by the user that can be used directly with an
      API, so the system must retrieve an API that can
@@
      ’recall_description’: ’WeatherForecastAPI(description:
      Retrieve weather forecast for a given location and date,
      input: location:string/city name;
      date:date/forecast date; output: weather:string/weather
      condition, temperature:float/forecast temperature)’}
@@
    _weather) based on the recall description above, after which the TWNM is invoked to acquire the corresponding call chains for these APIs (see Figure 10).
@@
In this round, the agent selects and executes the optimal API chain, with api-getweatherforecast_for_apjoy_weather_forecast designated as the target API.
ACTION: Tool Execution
@@
  The weather forecast for tomorrow in Beijing is partly
  cloudy with a temperature of 25C and humidity of 60%.
@@ 1. Core Requirements:
      Must explicitly contain initial parameters for leaf-
     node APIs
      Implicitly requires chained API calls from leaf to
     root node
      Root node API’s output directly resolves the user’s
     problem
@@ 2. Dependency Chain Rules:
      The output of every leaf-node API must be utilized in
     downstream
@@ 3. Parameter Constraints:
   - Root API output must contain realistic values matching
   its schema
@@ 5. Response Structure:
  "query": "<Real-world scenario requiring sequential API
  calls>",
@@
        "<param>": "<value explicitly stated in user query
        or previous API output>"
@@
        "data": {"<field>": "<realistic resolution to
        query>"}
@@ E.3.1 NaviAgent Inference Prompts
You are an intelligent API coordination system. Respond
strictly according to the following rules:
@@
   - Decompose compound requests into independent ordered
   sub-intents
      Sequential dependencies first, Must execute in
     declared order
@@
       - Create self-contained decision unit, action must
       implement full
@@
       - Each action’s analysis focuses only on its own
       intent
@@ 2. Match intent with API list:
        - Validate Observation in user input to confirm
        target API success:
@@
            a) Complete parameters -> action=call_api
            (execute based on 3.1 dependency resolution)
@@
       a) Can be completed via dependent APIs -> Execute
       Rule 3.1
       b) Use Retrieval APIs resolve parameter deficiencies
       in API
@@
    a) Check required parameters of target API, first call
    dependent APIs.
    b) For each missing parameter, select APIs from
    dependencies not marked
@@
    c) If an input parameter of an API is unavailable, use
    retrieve_api to
@@
  "intent": "1. <precisely describe the specific problem
  addressed by the current action>
             2. <extract data segments directly related to
             the subtask from user input>",
@@
        1.Explicitly state the specific decision-making sub-
        intent
@@
        "When action=retrieve_api: Use ’APIName(description:
        API functionality, input: param:type/description;
        output:
        param:type/description)’ format with only core
        parameters (e.g.,
        ’StockAPI(description: Query stock price by symbol,
        input: symbol:string/stock symbol; output:
        price:float/current price)’)",
  "answer": "When actionin[direct_answer,clarify_intent]:
  Natural language response (interrogative sentences
  required)"
@@
- JSON array items MUST be sorted by dependency_rank in
ascending order
@@
[OR]Target API must be executed successfully, as shown in
the status
@@ 1. Parameter names must strictly match API documentation
2. The ’answer’ field for clarify_intent must contain
question words
@@ 4. When action in [retrieve_api]:
    - The recall_description field serves exclusively as an
    API retrieval identifier from predefined repositories.
    - parameter descriptions must distinguish between input
    and output  parameters, retaining only essential
    parameters
    - Each recall_description can only recall one
    api,multiple APIs require
@@ 5. APIs absent from Candidate APIs MUST NOT be invented
6. When action=call_api is permitted only when candidate
APIs exist and the target_api is present in the candidate
APIs.
7. The "action" field must be strictly limited to one of the
following four predefined operation types: call_api,
direct_answer,retrieve_api or clarify_intent.